# Large dilatational hyperelasticity of glasses en route to cavitation failure

### 13 Mar 2026 v1

@@ Abstract
Materials deform elasto-plastically and fail under various loading conditions, typically quantified by the stress triaxiality, which is the ratio between the dilatational (hydrostatic) stress and the deviatoric (shear-like) one. We show that the elasto-plastic deformation of glasses approaching failure qualitatively differ for large and small stress triaxiality levels. Specifically, in the former limit, glasses reveal a strong hyperelastic (nonlinear elastic) response with minute plasticity, largely independently of the quenching rate across the glass transition. Yet, glassy disorder gives rise to significant elastic (reversible) nonaffine deformation, accompanied by the formation of micro-cavities. A small fraction of the latter is irreversible, i.e., survives unloading prior to the onset of failure, and may serve as nucleation sites for failure in the form of large-scale cavitation, upon which the glass loses a significant fraction of its load-bearing capacity. These results are contrasted with glass behavior in the limit of vanishing stress triaxiality and their universality across different glass formers is demonstrated. Finally, the implications of our findings for understanding glass deformation and failure under realistic stress conditions are discussed.
@@ I Introduction
Here, we show that in fact the elasto-plastic deformation of glasses approaching failure qualitatively differ for large and small stress triaxiality levels. By employing large-scale computer simulations and an array of theoretical and computational tools, we show that under large stress triaxiality, glasses reveal a strong hyperelastic (nonlinear elastic) response with minute plasticity, largely independently of the quench rate through the glass transition, which strongly affects the state of glassy disorder/stability. Consequently, dilatational elastic softening is accounted for by a first-principles zero-strain nonlinear elastic expansion, which is terminated by a large-scale cavitation instability.
@@ III Microscopic measures of nonaffinity and structural rearrangements
Figure 2: (a) The dimensionless measure of nonaffine deformation $\eta_{{}_{\rm n.a.}}$ vs. strain under both dilation and shear (as in Fig. 1 but with $N\!=\!10^{5}$), see text for details and discussion. (b) The corresponding accumulated number of structural rearrangements (instabilities).
@@ IV Macroscopic reversibility and irreversibility revealed by unloading
The results reveal striking differences between shear and dilation. Figure 3a shows that unloading dilatational states with $\epsilon\!<\!\epsilon_{\rm c}$ leads to a recovery of the macroscopic initial state, with no significant hysteresis and residual plastic strain. It indicates that minute dissipation is involved in the loading process up to $\epsilon_{\rm c}$, i.e., that the corresponding deformation is predominantly elastic/reversible. This means that the dilatational softening observed in Fig. 1b mainly corresponds to hyperelasticity, i.e., to nonlinear elasticity, not to plasticity. Likewise, the dilatational nonaffinity of Fig. 2a is expected to be predominantly elastic and the corresponding structural rearrangements quantified in Fig. 2b to be small and/or reversible upon unloading. Significant hysteresis and residual plastic strains emerge only after the large stress drop occurring at $\epsilon_{\rm c}$, as shown in Fig. 3a.
@@
Figure 3: (a-b) The same as Fig. 1a, but with unloading to zero strain added (the unloading points are marked by diamonds and guiding arrows are added). (c-d) The same as panels (a-b), but for a deeply annealed (slowly quenched) glass sample. See text for discussion.
@@ V Nonequilibrium thermal history effects
The structural rearrangements quantified in Fig. 2b do not give rise to significant plastic dissipation and stress relaxation. This implies that they are either small in magnitude or reversible upon unloading or both. Reversible structural rearrangements, i.e., saddle-node bifurcations that are elastically generated during loading, are documented for well annealed glasses deformed under shear [51, 52, 53, 54]. We are therefore interested in exploring the differences in the deformation and failure modes of glasses under vastly different stress triaxiality levels also as a function of the glass thermal history.
@@
These results point to yet another major difference between the shear and dilatational response of glasses, indicating that while the dilatational response is predominantly elastic prior to large-scale cavitation, largely independently of the glassy state of disorder (degree of thermal annealing), the macroscopic reversibility-irreversibility properties of the shear response strongly depend on it. The shear and dilatational response in Fig. 3c-d differ in two additional important respects: first, the physical (and spatial) origin of the large stress drops observed in the two cases is markedly different, i.e., it is large-scale cavitation under dilation (see Fig. 6a and SI Appendix), while it is large-scale shear-banding under shear [58, 59, 60]. Second, nonlinear elasticity under shear prior to shear-banding is very small, while it is very large under dilation, i.e., $K$ can soften up to $75\%$ over a dilatational strain of $3\!-\!4\%$ (see SI Appendix, where the deeply annealed counterpart plot of Fig. 1b is presented).
@@ VI Microscopic irreversibility
Next, we quantify the magnitude of structural rearrangements, whether reversible or irreversible, under shear and dilation. In Fig. 4c, we present the probability $p(s)$ of dimensionless stress drops $s$ (see legend) during structural rearrangements under shear and dilation for quickly quenched (high $T_{\rm p}$) samples up to $3\%$ strain (smaller than $\epsilon_{\rm c}$). It is observed that stress drops under dilation are more than an order of magnitude smaller than those under shear. Hence, not only structural rearrangements under dilation tend to be more reversible but they are also significantly less intense, outweighing their larger rate observed in Fig. 2a and consistently with the significantly lower dilatational nonaffinity observed in Fig. 2a. Overall, these results explain the main observation in Fig. 3 that dilatational deformation up to large-scale cavitation is macroscopically reversible/elastic, largely independently of the glass thermal history. They do not, in themselves, explain its nonlinearity, i.e., the large dilatational hyperelasticity.
@@ VIII Irreversible micro-cavities and the onset of large-scale cavitation
Figure 6: (a) Spatial distributions of micro-cavities (labeled by volume $V_{\rm cav}$, see right colorbar) and the magnitude of particle-level nonaffine displacements (in red) along the loading-unloading path of the dilated glass of Fig. 1a (see strain values $\epsilon$ and labeled arrows). The zoom-in inset shows an irreversible micro-cavity with its surrounding particles, featuring $V_{\rm cav}\!=\!11.3a_{0}^{3}$ with $a_{0}\!\equiv\![V(\epsilon\!=\!0)/N]^{1/3}$, also marked on the $\epsilon\!=\!\epsilon_{\rm c}^{-}$ snapshot. See text for discussion. (b) $p(V_{\rm cav}/a_{0}^{3}$) at $\epsilon\!=\!0$, $\epsilon\!=\!\epsilon_{\rm c}^{-}$ and the unloaded state (see legend), where the tail of the latter follows $\sim\!(V_{\rm cav})^{-3}$. The vertical dashed line denotes the lower threshold for micro-cavity detection in panel (a).
@@ X Summary and outlook
Our findings are directly relevant for realistic deformation and failure of glasses, which generically take place under finite stress triaxiality levels, in many cases relatively high ones [26, 27]. They can be tested in experiments that are specifically designed to achieve high stress triaxiality levels. Our results also give rise to a plethora of questions and future investigation directions, briefly mentioned here. First, as already alluded to above, one needs to develop a better understanding of the physical conditions for the formation of irreversible micro-cavities and of the instability criterion associated with large-scale cavitation. Such a criterion may also shed light on the quantitatively different behavior revealed in silica. Moreover, one needs to characterize the spatiotemporal evolution of the cavity, which requires going beyond the AQS limit, and to understand its size selection.

### 20 Mar 2026 v2

@@ Abstract
Materials deform elasto-plastically and fail under various loading conditions, typically quantified by the stress triaxiality, which is the ratio between the dilatational (hydrostatic) stress and the deviatoric (shear-like) one. We show that the elasto-plastic deformation of glasses approaching failure qualitatively differ for large and small stress triaxiality levels. Specifically, in the former limit, glasses reveal a strong hyperelastic (nonlinear elastic) response with minute plasticity, largely independently of the quenching rate across the glass transition. Yet, glassy disorder gives rise to significant elastic (reversible) nonaffine deformation, accompanied by the formation of micro-cavities. A small fraction of the latter is irreversible, i.e., survives unloading prior to the onset of failure, and may serve as nucleation sites for failure in the form of large-scale cavitation, involving a topological transition accompanied by the formation of an internal free surface, upon which the glass loses a significant fraction of its load-bearing capacity. These results are contrasted with glass behavior in the limit of vanishing stress triaxiality and their universality across different glass formers is demonstrated. Finally, the implications of our findings for understanding glass deformation and failure under realistic stress conditions are discussed.
@@ I Introduction
Here, we show that in fact the elasto-plastic deformation of glasses approaching failure qualitatively differ for large and small stress triaxiality levels. By employing large-scale computer simulations and an array of theoretical and computational tools, we show that under large stress triaxiality, glasses reveal a strong hyperelastic (nonlinear elastic) response with minute plasticity, largely independently of the quench rate through the glass transition, which strongly affects the state of glassy disorder/stability. Consequently, dilatational elastic softening is accounted for by a first-principles zero-strain nonlinear elastic expansion, which is terminated by a large-scale cavitation instability. The latter is a topological transition upon which an internal free surface is spontaneously formed.
@@ III Microscopic measures of nonaffinity and structural rearrangements
Figure 2: (a) The dimensionless measure of nonaffine deformation $\eta_{{}_{\rm n.a.}}$ vs. strain under both dilation and shear (as in Fig. 1 but prior to large-scale cavitation and with $N\!=\!10^{5}$), see text for details and discussion. (b) The corresponding accumulated number of structural rearrangements (instabilities).
@@ IV Macroscopic reversibility and irreversibility revealed by unloading
The results reveal striking differences between shear and dilation. Figure 3a shows that unloading dilatational states with $\epsilon\!<\!\epsilon_{\rm c}$ leads to a recovery of the macroscopic initial state, with no significant hysteresis and residual plastic strain. It indicates that minute dissipation is involved in the loading process up to $\epsilon_{\rm c}$, i.e., the corresponding deformation is predominantly elastic/reversible. This means that the dilatational softening observed in Fig. 1b mainly corresponds to hyperelasticity, i.e., to nonlinear elasticity, not to plasticity. Likewise, the dilatational nonaffinity of Fig. 2a is expected to be predominantly elastic and the corresponding structural rearrangements quantified in Fig. 2b to be small and/or reversible upon unloading. Significant hysteresis and residual plastic strains emerge only after the large stress drop occurring at $\epsilon_{\rm c}$, as shown in Fig. 3a.
@@
Figure 3: (a-b) The same as Fig. 1a, but with unloading to zero stress added (the unloading points are marked by diamonds and guiding arrows are added). (c-d) The same as panels (a-b), but for a deeply annealed (slowly quenched) glass sample. See text for discussion.
@@ V Nonequilibrium thermal history effects
The structural rearrangements under dilation quantified in Fig. 2b do not give rise to significant plastic dissipation and stress relaxation. This implies that they are either small in magnitude or reversible upon unloading or both. Reversible structural rearrangements, i.e., saddle-node bifurcations that are elastically generated during loading, are documented for well annealed glasses deformed under shear [51, 52, 53, 54]. We are therefore interested in exploring the differences in the deformation and failure modes of glasses under vastly different stress triaxiality levels also as a function of the glass thermal history.
@@
These results point to yet another major difference between the shear and dilatational response of glasses, indicating that while the dilatational response is predominantly elastic prior to large-scale cavitation, largely independently of the glassy state of disorder (degree of thermal annealing), the macroscopic reversibility-irreversibility properties of the shear response strongly depend on it. The shear and dilatational response in Fig. 3c-d differ in two additional important respects: first, the physical (and spatial) origin of the large stress drops observed in the two cases is markedly different, i.e., it is large-scale cavitation under dilation (see Fig. 6a and SI Appendix), while it is large-scale shear-banding under shear [58, 59, 60], having no topological signature. Second, nonlinear elasticity under shear prior to shear-banding is very small, while it is very large under dilation, i.e., $K$ can soften up to $75\%$ over a dilatational strain of $3\!-\!4\%$ (see SI Appendix, where the deeply annealed counterpart plot of Fig. 1b is presented).
@@ VI Microscopic irreversibility
Next, we quantify the magnitude of structural rearrangements, whether reversible or irreversible, under shear and dilation. In Fig. 4c, we present the probability $p(s)$ of dimensionless stress drops $s$ (see legend) during structural rearrangements under shear and dilation for quickly quenched (high $T_{\rm p}$) samples up to $3\%$ strain (smaller than $\epsilon_{\rm c}$). It is observed that stress drops under dilation are more than an order of magnitude smaller than those under shear. Hence, not only structural rearrangements under dilation tend to be more reversible but they are also significantly less intense, outweighing their larger rate observed in Fig. 2b and consistently with the significantly lower dilatational nonaffinity observed in Fig. 2a. Overall, these results explain the main observation in Fig. 3 that dilatational deformation up to large-scale cavitation is macroscopically reversible/elastic, largely independently of the glass thermal history. They do not, in themselves, explain its nonlinearity, i.e., the large dilatational hyperelasticity.
@@ VIII Irreversible micro-cavities and the onset of large-scale cavitation
Figure 6: (a) Spatial distributions of micro-cavities (labeled by volume $V_{\rm cav}$, see right colorbar) and the magnitude of particle-level nonaffine displacements (in red) along the loading-unloading path of the dilated glass of Fig. 1a (see strain values $\epsilon$ and labeled arrows). The zoom-in inset shows an irreversible micro-cavity with its surrounding particles, featuring $V_{\rm cav}\!=\!11.3a_{0}^{3}$ with $a_{0}\!\equiv\![V(\epsilon\!=\!0)/N]^{1/3}$, also marked on the $\epsilon\!=\!\epsilon_{\rm c}^{-}$ snapshot. See text for discussion. (b) $p(V_{\rm cav}/a_{0}^{3}$) at $\epsilon\!=\!0$, $\epsilon\!=\!\epsilon_{\rm c}^{-}$ and the unloaded state (see legend), where the tail of the $\epsilon\!=\!\epsilon_{\rm c}^{-}$ distribution follows $\sim\!(V_{\rm cav})^{-3}$. The vertical dashed line denotes the lower threshold for the micro-cavities shown in panel (a).
@@ X Summary and outlook
Our findings are directly relevant for realistic deformation and failure of glasses, which generically take place under finite stress triaxiality levels, in many cases relatively high ones [26, 27], and involve material decohesion. They can be tested in experiments that are specifically designed to achieve high stress triaxiality levels. Our results also give rise to a plethora of questions and future investigation directions, briefly mentioned here. First, as already alluded to above, one needs to develop a better understanding of the physical conditions for the formation of irreversible micro-cavities and of the instability criterion associated with large-scale cavitation. Such a criterion may also shed light on the quantitatively different behavior revealed in silica. Moreover, one needs to characterize the spatiotemporal evolution of the cavity, which requires going beyond the AQS limit, and to understand its size selection.